# Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu

### v1

@@ 4.2 Code Generation
Code generation involves the automatic creation of executable code from natural language specifications, facilitating a more intuitive interface for programming (Chen et al., 2021).
@@ 4.2.1 Parameter-Frozen Paradigm
Recent advancements in code generation have been significantly propelled by the development of LLMs, with studies showcasing their proficiency in generating code in a zero-shot manner. Code LLMs, trained on both code and natural language, have a robust and amazing zero-shot learning capability for programming tasks (Nijkamp et al., 2022; Roziere et al., 2023). Moreover, CodeT5+ enriches the landscape by proposing a flexible encoder-decoder architecture and a suite of pretraining objectives, leading to notable improvements (Wang et al., 2023i). These models collectively push the boundary of what is achievable in code generation, offering promising avenues for zero-shot learning.
@@
Code generation is being revolutionized by few-shot learning. This technique allows models to create precise code snippets by learning from just minimal examples Lu et al. (2021). Chen et al. (2021), Allal et al. (2023), Li et al. (2023f), Luo et al. (2023c) and Christopoulou et al. (2022) illustrate the efficacy of few-shot learning, demonstrating an adeptness at code generation that surpasses its predecessors. The development of smaller, yet powerful models (Li et al., 2023g; Guo et al., 2024), further highlights accessibility of few-shot code generation technologies, making them indispensable tools in the arsenal of modern developers.
@@ 4.2.2 Parameter-Tuning Paradigm
Full-parameter tuning represents a pivotal strategy in enhancing code generation models, allowing comprehensive model optimization. Specifically, CodeT series (Wang et al., 2021, 2023i) epitomize this approach by incorporating code-specific pre-training tasks and architecture flexibility, respectively, to excel in both code understanding and generation. CodeRL (Le et al., 2022) and PPOCoder (Shojaee et al., 2023) introduce deep reinforcement learning, leveraging compiler feedback and execution-based strategies for model refinement, whereas StepCoder (Shojaee et al., 2023) advances this further by employing reinforcement learning, curriculum learning and fine-grained optimization techniques. These models collectively demonstrate significant improvements across a spectrum of code-related tasks, embodying the evolution of AI-driven programming aids.
@@
PET emerges as a pivotal adaptation in code tasks, striking a balance between performance and computational efficiency (Weyssow et al., 2023). Studies (Ayupov and Chirkova, 2022; Zhuo et al., 2024) exploring adapters and LoRA showcase PET’s viability on code understanding and generation tasks, albeit with limitations in generative performance.
@@ 4.3 Machine Translation
Machine translation is a classical task that utilize computers to automatically translate the given information from one language to another, striving for accuracy and preserving the semantic essence of the original material (Bahdanau et al., 2014).
@@ 4.3.1 Parameter-Frozen Paradigm
In the realm of zero-shot learning, Zhu et al. (2023a) and Wei et al. (2023b) enhance LLMs’ multilingual performance through cross-lingual and multilingual instruction-tuning, significantly improving translation tasks. OpenBA contributes to the bilingual model space, demonstrating superior performance in Chinese-oriented tasks with a novel architecture (Li et al., 2023c). These advancements highlight the potential of LLMs in aligning language in zero-shot settings.
@@
In the exploration of few-shot learning for machine translation (MT), recent studies present innovative strategies to enhance the capabilities of LLMs Li et al. (2023a); Huang et al. (2024). Lu et al. (2023b) introduce Chain-of-Dictionary Prompting (CoD) to improve the MT of rare words by in-context-learning in low-resource languages. Raunak et al. (2023) investigate the impact of demonstration attributes on in-context learning, revealing the critical role of output text distribution in translation quality. Together, these works illustrate the significant potential of few-shot learning and in-context strategies in advancing the field of MT with LLMs.
@@ 4.3.2 Parameter-Tuning Paradigm
Full-parameter tuning in machine translation with LLMs represents a frontier for enhancing translation accuracy and adaptability Xu et al. (2023b). Iyer et al. (2023) demonstrate the potential of LLMs in disambiguating polysemous words through in-context learning and fine-tuning on ambiguous datasets, achieving superior performance in multiple languages. Moslem et al. (2023) and Wu et al. (2024) focus on exploring fine-tuning methods that enhance real-time and context-aware translation capabilities. Xu et al. (2024) propose Contrastive Preference Optimization (CPO) to refine translation quality further, pushing LLMs towards better performance. These studies reveal the efficacy and necessity of fine-tuning approaches in realizing the full potential of LLMs for complex machine translation tasks.
@@
PET is emerging as a transformative approach for integrating LLMs into machine translation (MT), balancing performance and efficiency. Ustun and Stickland (2022) empirically assess PET’s efficacy across different languages and model sizes, highlighting adapters’ effectiveness with adequate parameter budgets. Alves et al. (2023) optimize the finetuning process with adapters, striking a balance between few-shot learning and finetuning efficiency. These studies collectively underline PET’s potential to revolutionize MT by making LLMs more adaptable and resource-efficient.
@@ 4.4 Mathematical Reasoning
Mathematical reasoning tasks in NLP involve the use of NLP techniques to understand information from mathematical text, perform logical reasoning, and generate answers Lu et al. (2023e).
@@ 4.4.1 Parameter-Frozen Paradigm
Mathematics serves as a testbed to investigate the reasoning capabilities of LLMs OpenAI (2023); Touvron et al. (2023). The vanilla prompting method asks LLMs to directly arrive at the final answer to a given mathematical problem. It is very challenging and the reasoning process is not transparent to humans. To address it,  Kojima et al. (2022) develop a zero-shot chain-of-thought technique, which utilizes the simple prompt “Let’s think step by step” to elicit mathematical reasoning in LLMs. By doing this, the LLM can break down the problem into smaller, easier-to-solve pieces before arriving at a final answer. Further, Wang et al. (2023g) propose a new decoding strategy, called self-consistency. This approach integrates a series of prompting results to boost the mathematical performance.
@@
Recent studies explore constructing more suitable exemplars for LLMs to improve mathematical reasoning.  Wei et al. (2022c) introduce chain-of-thought prompting, which presents a few chain-of-thought demonstrations to teach LLMs to think step by step. However, manually constructing the demonstrations in few-shot learning is time- and labor-consuming. To solve this problem,  Zhang et al. (2022b) and  Lu et al. (2023d) propose to select in-context examples automatically. Even given detailed examples, it is still hard for LLMs to calculate the numbers precisely. To address this issue, PAL Gao et al. (2023b) directly generates programs as intermediate reasoning steps. These programs are then executed using a runtime environment, like a Python interpreter, to find the better and robust solution.
@@ 4.4.2 Parameter-Tuning Paradigm
Full-parameter tuning is a common way to specify LLMs’ behaviors on mathematical reasoning tasks.  Luo et al. (2023a) apply their proposed Reinforcement Learning from Evol-Instruct Feedback (RLEIF) method to the domain of math to improve the mathematical reasoning abilities of LLMs.  Yue et al. (2023) introduce the MathInstruct dataset to enhance the general math problem-solving ability of LLMs through in-domain instruction tuning.  Ho et al. (2023) teach the small language models to perform mathematical reasoning by distilling the generated intermediate rationales by large language models.  Schick et al. (2023) present ToolFormer, which can use the calculator to perform simple numeric calculations when solving math problems.
@@
Fine-tuning LLMs with full parameter updates incurs significant memory overhead, limiting accessibility for many users. Parameter-efficient tuning techniques, such as LoRA Hu et al. (2022a), offer a promising alternative. Additionally, Hu et al. (2023b) propose a user-friendly framework for integrating various adapters into LLMs, enabling them to tackle tasks like mathematical reasoning.
@@ 5 Future Work and New Frontier
In this section, as shown in Figure 4, we highlight some new frontiers, hoping to spur more breakthroughs in the future.
@@ 5.1 Multilingual LLMs for NLP
Despite the significant success of LLMs in English NLP tasks, there are over 7,000 languages worldwide. How to extend the success of English-centric LLMs to NLP tasks in other languages is an important research question Qin et al. (2024). Inspired by this, recent research has increasingly focused on using multilingual LLMs to solve NLP tasks in multilingual scenarios Xue et al. (2021); Workshop et al. (2022); Shi et al. (2022); Qin et al. (2023a); Winata et al. (2023).
Two main challenges in this direction are as follows: (1) Enhancing Low-Resource Language Performance: Due to poor performance in low-resource languages, how to build universal multilingual LLMs that achieve promising performance in NLP tasks across languages is a direction worth exploring. (2) Improving Cross-lingual Alignment: The key to multilingual LLMs is improving the alignment between English and other languages. Effectively achieving cross-lingual alignment in cross-lingual NLP tasks is a challenge.
@@ 5.2 Multi-modal LLMs for NLP
The current LLMs achieve excellent performance in text modality. However, integrating more modalities is one of the key ways to achieve artificial general intelligence (AGI). Therefore, a lot of work has begun to explore multi-modal LLMs for multi-modal NLP tasks Lu et al. (2022, 2023c); Yang et al. (2023a, b); Zhang et al. (2023l).
The primary challenges in this field are: (1) Complex Multi-modal Reasoning: Currently, most multi-modal LLMs focus on simple multi-modal reasoning, like recognition Wang et al. (2023e); Liu et al. (2023a), while neglecting complex multi-modal reansoning Yang et al. (2023b); Lu et al. (2023c). Therefore, how to effectively explore complex multi-modal reasoning for NLP is a crucial topic. (2) Effective Multi-modal Interaction: Existing methods often simply focus on adding direct multi-modal projection or prompting to LLM for bridge multi-modality gap Wang et al. (2023e); Liu et al. (2023a); Wu et al. (2023b); Mitra et al. (2023). Crafting a more effective multi-modal interaction mechanism in multi-modal LLMs to solve NLP tasks is an essential problem.
@@ 5.3 Tool-usage in LLMs for NLP
While LLMs have shown success in NLP tasks, they can still face challenges when applied in real-world scenarios Qin et al. (2023b). Therefore, a lot of work explores utilizing LLMs as central controllers to enable the usage or construction of tools and agents to solve more practical NLP tasks Shinn et al. (2023); Wang et al. (2023c); Zhu et al. (2023b); Hu et al. (2023a).
@@ 5.4 X-of-thought in LLMs for NLP
When LLMs solve complex NLP problems, they often cannot directly give correct answers and require complex thinking. Therefore, some works adapt X-of-thought (XoT) for advanced logical reasoning. XoT primarily aims to refine logical processing for better NLP task solution Kojima et al. (2022); Zhang et al. (2022b); Qin et al. (2023a); Yao et al. (2023); Chen et al. (2022); Lei et al. (2023).

### v2

@@ 4.2 Code Generation
Code generation involves the automatic creation of executable code from natural language specifications, facilitating a more intuitive interface for programming [98].
@@ 4.2.1 Parameter-Frozen Paradigm
Recent advancements in code generation have been significantly propelled by the development of LLMs, with studies showcasing their proficiency in generating code in a zero-shot manner. Code LLMs, trained on both code and natural language, have a robust and amazing zero-shot learning capability for programming tasks [99, 106]. Moreover, CodeT5+ enriches the landscape by proposing a flexible encoder-decoder architecture and a suite of pretraining objectives, leading to notable improvements [155]. These models collectively push the boundary of what is achievable in code generation, offering promising avenues for zero-shot learning. Recent releases of code-specific LLMs, such as CodeGemma [211] and Qwen2.5-Coder [212], further advance the field of LLM-based code generation, delivering superior benchmark performance. Additionally, Seed-Coder [213] introduces a model-centric data curation pipeline, while Ling-Coder-Lite [214] leverages a Mixture-of-Experts architecture to balance efficiency and performance, marking state-of-the-art progress in open-source code generation LLMs.
@@
Code generation is being revolutionized by few-shot learning. This technique allows models to create precise code snippets by learning from just minimal examples [215]. Chen et al. [98], Allal et al. [102], Li et al. [103], Luo et al. [101], and Christopoulou et al. [100] illustrate the efficacy of few-shot learning, demonstrating an adeptness at code generation that surpasses its predecessors. The development of smaller, yet powerful models [104, 105] further highlights the accessibility of few-shot code generation technologies, making them indispensable tools in the arsenal of modern developers. Importantly, most modern LLMs for code generation, including Code Llama [106], Seed-Coder [213], Qwen2.5-Coder [212], and CodeGemma [211], provide both base and instruct variants, enabling flexible few-shot learning execution across diverse programming tasks.
@@ 4.2.2 Parameter-Tuning Paradigm
Full-parameter tuning represents a pivotal strategy in enhancing code generation models, allowing comprehensive model optimization. Specifically, CodeT series [154, 155] epitomize this approach by incorporating code-specific pre-training tasks and architecture flexibility, respectively, to excel in both code understanding and generation. CodeRL [156] and PPOCoder [157] introduce deep reinforcement learning, leveraging compiler feedback and execution-based strategies for model refinement, whereas StepCoder [157] advances this further by employing reinforcement learning, curriculum learning, and fine-grained optimization techniques. These models collectively demonstrate significant improvements across a spectrum of code-related tasks, embodying the evolution of AI-driven programming aids. Emerging work such as PRLCoder [216] leverages process-supervised reinforcement learning, Focused-DPO [217] enhances preference optimization on error-prone points, and ACECoder [218] applies automated test-case synthesis to refine reward models. Furthermore, SWE-RL [219] expands reinforcement learning into real-world software engineering, significantly advancing the reasoning capacities of LLMs. Reinforcement learning thus demonstrates strong potential for training code LLMs and warrants further exploration.
@@
PET emerges as a pivotal adaptation in code tasks, striking a balance between performance and computational efficiency [160]. Studies [158, 159] exploring adapters and LoRA showcase PET’s viability on code understanding and generation tasks, albeit with limitations in performance. Recent investigations, such as Storhaug and Li [220], demonstrate that PEFT methods can rival full fine-tuning for unit test generation, reducing resource demands. Additionally, Zhang et al. [221] provide a comprehensive evaluation of PEFT on method-level code smell detection, revealing that small models often perform competitively, reinforcing the scalability and cost-effectiveness of PET techniques for specialized software engineering tasks.
@@ 4.3 Machine Translation
Machine translation is a classical task that utilizes computers to automatically translate the given information from one language to another, striving for accuracy and preserving the semantic essence of the original material [222]. Recent work [223] revisits key challenges in neural machine translation (NMT), highlighting how LLMs address issues such as long sentence translation and reduced parallel data reliance while facing new challenges like inference efficiency and low-resource language translation.
@@ 4.3.1 Parameter-Frozen Paradigm
In the realm of zero-shot learning, Zhu et al. [109] and Wei et al. [108] enhance LLMs’ multilingual performance through cross-lingual and multilingual instruction-tuning, significantly improving translation tasks. OpenBA contributes to the bilingual model space, demonstrating superior performance in Chinese-oriented tasks with a novel architecture [111]. These advancements highlight the potential of LLMs in aligning language in zero-shot settings.
@@
In the exploration of few-shot learning for machine translation (MT), recent studies present innovative strategies to enhance the capabilities of LLMs [110, 224]. Lu et al. [114] introduce Chain-of-Dictionary Prompting (CoD) to improve the MT of rare words by in-context learning in low-resource languages. Raunak et al. [113] investigate the impact of demonstration attributes on in-context learning, revealing the critical role of output text distribution in translation quality. Zhu et al. [225] propose a robust multi-view approach for selecting fine-grained demonstrations, effectively reducing noise in in-context learning and significantly improving domain adaptation. Together, these works illustrate the significant potential of few-shot learning in advancing the field of MT with LLMs.
@@ 4.3.2 Parameter-Tuning Paradigm
Full-parameter tuning in machine translation with LLMs represents a frontier for enhancing translation accuracy and adaptability [161]. Iyer et al. [163] demonstrate the potential of LLMs in disambiguating polysemous words through in-context learning and fine-tuning on ambiguous datasets, achieving superior performance in multiple languages. Moslem et al. [164] and Wu et al. [167] focus on exploring fine-tuning methods that enhance real-time and context-aware translation capabilities. Xu et al. [162] propose Contrastive Preference Optimization (CPO) to refine translation quality further, pushing LLMs towards better performance. Feng et al. [226] introduce MT-R1-Zero, applying reinforcement learning frameworks to MT without supervised fine-tuning, achieving competitive results on multilingual benchmarks and offering insights into emergent reasoning patterns. Feng et al. [227] present MT-Ladder, a cost-effective hierarchical fine-tuning framework that boosts general-purpose LLMs’ translation performance to match state-of-the-art models. These studies reveal the efficacy and necessity of fine-tuning approaches, and point toward reinforcement learning as a promising future direction for advancing machine translation by leveraging LLMs’ emergent reasoning and adaptability.
@@
PET is emerging as a transformative approach for integrating LLMs into machine translation (MT), balancing performance and efficiency. Ustun and Stickland [165] empirically assess PET’s efficacy across different languages and model sizes, highlighting adapters’ effectiveness with adequate parameter budgets. Alves et al. [112] optimize the fine-tuning process with adapters, striking a balance between few-shot learning and fine-tuning efficiency. Recent work further demonstrates PET’s scalability and robustness in multilingual and domain-specific tasks, confirming its potential to make LLMs more adaptable and resource-efficient while maintaining competitive performance. These studies collectively underline PET’s promise to revolutionize MT by offering scalable and cost-effective solutions.
@@ 4.4 Mathematical Reasoning
Mathematical reasoning tasks in NLP involve the use of NLP techniques to understand information from mathematical text, perform logical reasoning processes, and ultimately generate accurate answers to mathematical questions [228, 229].
@@ 4.4.1 Parameter-Frozen Paradigm
Mathematics serves as a testbed to investigate the reasoning capabilities of LLMs [230, 14]. The vanilla prompting method asks LLMs to directly arrive at the final answer to a given mathematical problem. It is very challenging and the reasoning process is not transparent to humans. To address it, Kojima et al. [31] develop a zero-shot chain-of-thought technique, which utilizes the simple prompt “Let’s think step by step” to elicit mathematical reasoning in LLMs. By doing this, the LLM can break down the problem into smaller, easier-to-solve pieces before arriving at a final answer. Further, Wang et al. [116] propose a new decoding strategy, called self-consistency. This approach integrates a series of prompting results to boost the performance. Tang et al. [231] propose an automatically enhanced zero-shot prompting strategy that adjusts the prompts through model retrieval to improve the performance of LLMs on mathematical reasoning tasks. Moreover, Yuksekgonul et al. [232] and Peng et al. [233] employ reflection-based, iterative prompting strategies to improve zero-shot mathematical reasoning accuracy.
@@
Recent studies explore constructing more suitable exemplars for LLMs to improve mathematical reasoning. Wei et al. [16] introduce chain-of-thought prompting, using a few demonstrations to guide LLMs through step-by-step reasoning. However, creating these examples by hand is laborious, so Zhang et al. [115] and Lu et al. [117] propose methods to select in-context examples automatically. To improve numerical precision, PAL [118] generates and executes intermediate program steps in a runtime environment. Building on this idea, Das et al. [119] present MathSensei, a tool-augmented LLM that integrates web search, code execution, and symbolic solving, showing greater gains on harder problems. Liu et al. [177] propose XoT, a unified framework that dynamically switches among diverse prompting methods for better math reasoning. To probe consistency, Yu et al. [178] use symbolic programs to reveal that LLMs often rely on brittle reasoning despite strong static performance. More recently, Ranaldi et al. [179] introduce QuaSAR, which blends natural language with selective formalization to enhance chain-of-thought robustness without full symbolic translation. Moreover, Zhang et al. [234] enhance the mathematical capabilities of LLMs by improving their single-step reasoning in the context of fine-grained in-context learning.
@@ 4.4.2 Parameter-Tuning Paradigm
Full-parameter tuning is a standard method for guiding LLMs in mathematical reasoning tasks [181]. Several studies have improved general math-solving ability by creating high-quality instruction-tuning datasets, from web-curated collections [169], advanced LLM distillation [170], and self-generated samples [235, 181]. Moreover, Schick et al. [171] introduce ToolFormer, which leverages a calculator for numeric operations. Chen et al. [176] propose perturbing token-level chain-of-thought during fine-tuning, improving accuracy without external labels. Yu et al. [236] develop Chain-of-Reasoning, which integrates natural-language, algorithmic, and symbolic reasoning to boost benchmarks. Beyond supervised tuning, reinforcement learning has also shown promise. Luo et al. [168] apply RLEIF to enhance math reasoning; Luo et al. [175] propose OmegaPRM, an MCTS-based method for training reward models on MATH500 and GSM8K without human oversight; Shao et al. [174] train DeepSeekMath 7B on 120 B tokens using web data and GRPO; and Qian et al. [237] introduce ToolRL, examining tool selection and reward design in RL-based fine-tuning.
@@
Fine-tuning LLMs with full parameter updates incurs significant memory overhead, limiting accessibility for many users. Parameter-efficient tuning techniques, such as LoRA [37], offer a promising alternative. Additionally, Hu et al. [172] propose a user-friendly framework for integrating various adapters into LLMs, enabling them to tackle tasks like mathematical reasoning. SPHERE [238] introduces a self‑evolving data‑generation pipeline leveraging LoRA to enhance the performance of small‑scale language models on mathematical reasoning tasks through the self‑generation, refinement, and diversification of reasoning chains. Prottasha et al. [239] present Semantic Knowledge Tuning (SK‑Tuning), which employs semantically meaningful vocabulary in lieu of random tokens for prompt and prefix tuning, thereby boosting LLM performance on mathematical reasoning tasks. Srivastava et al. [180] propose DTE, a ground truth-free training framework using multi-agent debates and a Reflect-Critique-Refine strategy to enhance LLM reasoning, achieving notable accuracy gains and strong cross-domain generalization. Further, Alazraki and Rei [240] introduce a meta‑reasoning‑based tool selection framework, a two‑stage system first performs meta‑reasoning over the given task and then leverages a custom, fine‑tuned language‑modeling head to generate candidate tools, thereby substantially improving mathematical reasoning performance.
@@ 5 Future Work and New Frontier
In this section, as shown in Figure 4, we highlight some new frontiers, aiming to inspire further innovations and groundbreaking advancements in the near future.
@@ 5.1 Multilingual LLMs for NLP
Despite the significant success of LLMs in English NLP tasks, there are over 7,000 languages worldwide. How to extend the success of English-centric LLMs to NLP tasks in other languages is an important research question [241, 242, 243, 244, 245]. Inspired by this, Researchers have made efforts to enhance the multilingual LLM through parameter-tuning strategies, including multilingual pretraining [246, 247, 248, 249], supervised fine-tuning [248, 249, 250], and reinforcement learning [251]. Other studies focus on cross-lingual alignment via prompting, using few-shot approaches [252, 253, 254, 255] and zero-shot instructions [256, 257] to enhance alignment.
Two main challenges in this direction are as follows: (1) Enhancing Low-Resource Language Performance: Due to poor performance in low-resource languages, how to build universal multilingual LLMs that achieve promising performance in NLP tasks across languages is a direction worth exploring. (2) Improving Cross-lingual Alignment: The key to multilingual LLMs is improving the alignment between English and other languages. Effectively achieving this alignment is critical for ensuring optimal performance in cross-lingual NLP tasks, making it a challenging yet essential area for advancement.
@@ 5.2 Multi-modal LLMs for NLP
The current LLMs achieve excellent performance in text modality. However, integrating modalities is one of the key ways to achieve AGI [258, 259, 260, 261]. Therefore, a lot of work has begun to explore multi-modal LLMs for multi-modal NLP tasks [262, 263, 264, 265, 266, 267, 268].
The primary challenges in this field are: (1) Complex Multi-modal Reasoning: Currently, most multi-modal LLMs focus on simple multi-modal reasoning, like recognition [269, 270], while neglecting complex multi-modal reansoning [271, 272, 273]. Therefore, how to effectively explore complex multi-modal reasoning for NLP is a crucial topic [274, 275, 259, 276]. (2) Effective Multi-modal Interaction: Existing methods often simply focus on adding direct multi-modal projection or prompting to LLM for bridge multi-modality gap [269, 270, 277, 278, 279]. Crafting a more effective multi-modal interaction mechanism in the inference process of multi-modal LLMs to solve NLP tasks is an essential problem.
@@ 5.3 Tool-usage in LLMs for NLP
While LLMs have shown success in NLP tasks, they can still face challenges when applied in real-world scenarios [280, 281]. Therefore, a lot of work focuses on exploring utilizing LLMs as central controllers to enable the usage or construction of tools and agents to solve practical NLP tasks [282, 283, 284, 285, 286, 287].
@@ 5.4 X-of-thought in LLMs for NLP
When LLMs solve complex NLP problems, they often cannot directly give correct answers and require complex thinking. Therefore, some works adapt X-of-thought (XoT) for advanced logical reasoning. XoT primarily focuses on refining the model’s ability to process and reason through complex logic, ultimately aiming to improve the overall performance and accuracy in solving challenging NLP tasks [31, 115, 256, 288, 289, 290, 291].